# Neural Conditional Transport Maps

### v1

@@ 3.2 Neural Optimal Transport
Transport & Critic: The transport map $T:\mathcal{X}\times\mathcal{Z}\times\mathcal{C}\rightarrow\mathcal{Y}$ is implemented as a neural network with encoder-decoder structure. For the common case where $\mathcal{X}=\mathcal{Y}=\mathbb{R}^{n}$, the encoder maps $\mathbb{R}^{n+s}$ into $\mathbb{R}^{d}$, a conditioning module transforms $\mathbb{R}^{d}\times\mathcal{C}$ into $\mathbb{R}^{d^{\prime}}$, and the decoder maps $\mathbb{R}^{d^{\prime}}$ into $\mathbb{R}^{n}$. The complete transport map is: $T(x,z,c)=\text{Decoder}_{T}(\text{Conditioning}_{T}(\text{Encoder}_{T}([x,z]),c))$ where $[x,z]\in\mathbb{R}^{n+s}$ denotes the concatenation of input $x$ and noise $z$, $d$ is the latent dimension, and $d^{\prime}$ is the conditioned latent dimension. The critic function $f:\mathcal{Y}\times\mathcal{C}\rightarrow\mathbb{R}$ follows the same structure without noise input: $f(y,c)=\text{Decoder}_{f}(\text{Conditioning}_{f}(\text{Encoder}_{f}(y),c))$
Layer Design: Unlike CondOT (Bunne et al., 2022), which uses ICNNs due to convexity assumptions, our architecture supports more flexible designs. We use residual blocks He et al. (2016) for both $T$ and $f$, along with layer normalization Ba et al. (2016) and orthogonal initialization Saxe et al. (2014). Each block consists of $\text{ResBlock}(h)=h+\alpha\cdot\mathcal{F}(h)$, where $\alpha$ is a learnable scaling parameter and $\mathcal{F}$ represents a composite function that may include normalization, non-linear activations, and appropriate transformations for the data type.
Encoder-Decoder Design: The encoder produces a sequence of hidden states $\{h_{1},h_{2},...,h_{L}\}$ where $h_{L}\in\mathbb{R}^{d}$ is used for conditioning. The decoder takes the conditioned representation and maps to the output space. Both modules can be instantiated with different layer types depending on the data modality: $\text{Encoder}(x)=h_{L}$ where $h_{i}=\text{Layer}_{i}(h_{i-1})$ and $h_{0}=x$. Here, $\text{Layer}_{i}$ can be any differentiable layer. The encoder learns condition-invariant features from the source distribution $\mathbb{P}$, while the decoder applies condition-specific transformations.
@@
1:Input: Distributions $\mathbb{P}$, $\mathbb{Q}$, $\mathbb{S}$ accessible by samples; Transport network $T_{\theta}\colon\mathbb{R}^{P}\times\mathbb{R}^{S}\times\mathcal{C}\rightarrow\mathbb{R}^{Q}$; Critic network $f_{\omega}\colon\mathbb{R}^{Q}\times\mathcal{C}\rightarrow\mathbb{R}$;
2:            Cost $\mathcal{L}\colon\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$; Conditioning distribution $\mathcal{C}$; Maximum steps $N$, Transport iterations per step $K_{T}$
3:Pre-training: Initialize $T_{\theta}$ and $f_{\omega}$ using objectives in Eq. (3)-(7)
4:Output: Learned conditional transport map $T_{\theta}$
5:for $t=1,2,\ldots,N$ do
6:    Sample conditioning $c\sim\mathcal{C}$
7:    Sample batches $Y\sim\mathbb{Q}$, $X\sim\mathbb{P}$; for each $x\in X$ sample $Z_{x}\sim\mathbb{S}$
8:    $\mathcal{L}_{f}\leftarrow\frac{1}{|X|}\sum_{x\in X}\frac{1}{|Z_{x}|}\sum_{z\in Z_{x}}f_{\omega}(T_{\theta}(x,z,c),c)-\frac{1}{|Y|}\sum_{y\in Y}f_{\omega}(y,c)$
9:    Update $\omega$ using $\frac{\partial\mathcal{L}_{f}}{\partial\omega}$ (gradient ascent)
10:    for $k_{T}=1,2,\ldots,K_{T}$ do
11:       Sample conditioning $c\sim\mathcal{C}$
12:       Sample batch $\tilde{X}\sim\mathbb{P}$; for each $x\in\tilde{X}$ sample $\tilde{Z}_{x}\sim\mathbb{S}$
13:       $\mathcal{L}_{T}\leftarrow\frac{1}{|\tilde{X}|}\sum_{x\in\tilde{X}}[\mathcal{L}(x,T_{\theta}(x,\tilde{Z}_{x},c))-\frac{1}{|\tilde{Z}_{x}|}\sum_{z\in\tilde{Z}_{x}}f_{\omega}(T_{\theta}(x,z,c),c)]$
14:       Update $\theta$ using $\frac{\partial\mathcal{L}_{T}}{\partial\theta}$
@@ 3.3 Conditioning mechanism
Our framework supports conditioning on discrete and continuous variables, with the capability to enable either or both types of conditioning depending on the application. The conditioning module transforms these inputs into a unified representation that modulates the transport map.
@@
Unified Conditioning. The conditioning module flexibly handles discrete variables, continuous variables, or both. When both types are present, their representations are concatenated, then processed to produce the final conditioning vector $\mathbf{c}=\text{MLP}([\mathbf{e}_{\text{discrete}},\text{PE}(p_{\text{continuous}})])$, where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation and MLP is a multi-layer perceptron with SiLU Ramachandran et al. (2017) activations. The resulting vector $\mathbf{c}\in\mathbb{R}^{d_{c}}$ is used to modulate the transport map in the latent space.
Flexible Configuration. Both discrete and continuous conditioning are optional and can be independently enabled or disabled. The module requires at least one type of conditioning to be active. This flexibility allows our framework to adapt to various application domains—from purely categorical problems to continuous spatiotemporal transport tasks. Unlike CondOT (Bunne et al., 2022), we use different conditioning embeddings for T and f, as this showed better performance.
@@ 3.4 Conditioning modules variants
Hypernetwork Conditioning The hypernetwork generates the final layer weights and biases of the encoder based on the conditioning. Given the encoder output $\mathbf{h}\in\mathbb{R}^{d}$ and conditioning $\mathbf{c}$, the hypernetwork $\mathcal{H}$ is a shallow MLP that generates parameters: $[\mathbf{W},\mathbf{b}]=\mathcal{H}(\mathbf{c}),\text{where }\mathbf{W}\in\mathbb{R}^{d\times n},\mathbf{b}\in\mathbb{R}^{n}$. The final output is then computed as: $\mathbf{y}=\mathbf{h}\mathbf{W}+\mathbf{b}$. Unlike in feature modulation, generating weights allows fundamentally different transformations per condition, essential for optimal transport where different conditions require distinct mapping strategies, providing the expressiveness of separate networks per condition without its increased computational cost.
Alternative Conditioning Mechanisms We evaluated several conditioning strategies beyond our hypernetwork approach. The simplest baseline is Concatenation, which directly combines feature vectors with the condition. More sophisticated approaches include Cross-Attention (Vaswani et al., 2017), which uses attention mechanisms to modulate features based on conditions, and Feature-wise Linear Modulation (FiLM) (Perez et al., 2018), which applies learnable transformations to features. We also tested normalization-based methods like Adaptive Instance Normalization (Huang and Belongie, 2017) and Conditional Layer Normalization (Su et al., 2021), along with attention-inspired techniques such as Squeeze-and-Excite (Hu et al., 2018) and Feature-wise Affine Normalization (FAN) (Zhou et al., 2021). Although these alternatives showed promise in specific scenarios, our hypernetwork approach consistently demonstrated superior performance across all benchmarks. Notably, our lightweight hypernetwork implementation proved to be computationally efficient both in training time and parameter count, making it a sound choice even for low-budget scenarios. We ablate these components on Section 4 and provide more details in the supplementary material.
@@ 3.5 Training Procedure
During training, the transport map $T$ and critic function $f$ must maintain adversarial balance while adapting to diverse conditioning values. This creates optimization challenges. We address them through a two-phase approach: pre-training for stable initialization followed by minimax optimization.
Pre-training: Randomly initialized networks may implement highly non-linear transformations far from identity, leading to unstable optimization. Motivated by this observation, we introduce a lightweight pre-training that establishes favorable initial conditions for both networks. We pre-train $T$ to approximate the identity mapping: $\mathcal{L}_{T}^{\text{pre}}=\mathbb{E}_{x\sim\mathbb{P},z\sim\mathbb{S},c\sim\mathcal{C}}[\|T(x,z,c)-x\|_{2}^{2}]$. This initialization ensures that, early in training, the transport map preserves input structure. Besides, we pre-train $f$ with a multi-objective loss: $\mathcal{L}_{f}^{\text{pre}}=\lambda_{\text{smooth}}\mathcal{L}_{\text{smooth}}+\lambda_{\text{transport}}\mathcal{L}_{\text{transport}}+\lambda_{\text{mag}}\mathcal{L}_{\text{mag}}$. First, the smoothness term prevents sharp discontinuities that lead to gradient explosion during adversarial training:$\mathcal{L}_{\text{smooth}}=\mathbb{E}_{x\sim\mathbb{P},c\sim\mathcal{C}}[\|f(x+\epsilon,c)-f(x,c)\|_{2}^{2}]$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. Second, the transport term maintains the OT objective:$\mathcal{L}_{\text{transport}}=\mathbb{E}_{x\sim\mathbb{P},z\sim\mathbb{S},c\sim\mathcal{C}}[f(T(x,z,c),c)]-\mathbb{E}_{y\sim\mathbb{Q},c\sim\mathcal{C}}[f(y,c)]$. This helps $f$ learn meaningful discrimination between transported and target samples from initialization. Finally, a magnitude control term prevents unbounded growth, enhancing stability:$\mathcal{L}_{\text{mag}}=\mathbb{E}_{y\sim\mathbb{Q},c\sim\mathcal{C}}[(|f(y,c)|-1)^{2}]$.
@@
Optimization: Following pre-training, we employ an alternating training that reflects the minimax structure of optimal transport. We detail our approach in Algorithm 1. Following Korotin et al. (2023), we perform $K_{T}>1$ updates for the transport map per critic update. This asymmetry addresses the imbalance in learning complexity—the transport map must learn condition-dependent transformations while maintaining transport constraints, whereas the critic primarily serves as a measure of transport quality. We find $K_{T}\in\{4,6\}$ provides optimal balance between training stability and efficiency.
@@
Conditions Sampling: The performance of this learning procedure depends on how we sample from the conditioning space $\mathcal{C}$ during training. For discrete variables, we use uniform sampling across all categories. For continuous variables, we employ Beta distribution sampling $\text{Beta}(\alpha,\beta)$ with symmetric parameters ($\alpha=\beta=0.95$), which slightly oversamples values around the min and max values in the training datasets. This strategy prevents underfitting in boundary conditions.

### v2

@@ 3.2 Neural Optimal Transport
Transport & Critic: The transport map $T:\mathcal{X}\times\mathcal{Z}\times\mathcal{C}\rightarrow\mathcal{Y}$ is implemented as a neural network with encoder-decoder structure. For the common case where $\mathcal{X}=\mathcal{Y}=\mathbb{R}^{n}$, the encoder maps $\mathbb{R}^{n+s}$ into $\mathbb{R}^{d}$, a conditioning module transforms $\mathbb{R}^{d}\times\mathcal{C}$ into $\mathbb{R}^{d^{\prime}}$, and the decoder maps $\mathbb{R}^{d^{\prime}}$ into $\mathbb{R}^{n}$. The complete transport map is: $T(x,z,c)=\text{Decoder}_{T}(\text{Conditioning}_{T}(\text{Encoder}_{T}([x,z]),c))$ where $x,z]\in\mathbb{R}^{n+s}$ denotes the concatenation of input $x$ and noise $z$, $d$ is the latent dimension, and $d^{\prime}$ is the conditioned latent dimension. The critic function $f:\mathcal{Y}\times\mathcal{C}\rightarrow\mathbb{R}$ follows the same structure without noise input: $f(y,c)=\text{Decoder}_{f}(\text{Conditioning}_{f}(\text{Encoder}_{f}(y),c))$
Layer Design: Unlike CondOT (Bunne et al., 2022), which uses ICNNs, our unconstrained architecture supports flexible designs with standard optimization procedures. We use residual blocks He et al. (2016) for both $T$ and $f$, along with layer normalization Ba et al. (2016) and orthogonal initialization Saxe et al. (2014). Each block consists of $\text{ResBlock}(h)=h+\alpha\cdot\mathcal{F}(h)$, where $\alpha$ is a learnable scaling parameter and $\mathcal{F}$ represents a composite function that may include normalization, non-linear activations, and appropriate transformations for the data type.
Encoder-Decoder Design: The encoder produces a sequence of hidden states $\{h_{1},h_{2},...,h_{L}\}$ where $h_{L}\in\mathbb{R}^{d}$ is used for conditioning. The decoder takes the conditioned representation and maps to the output space. Both modules can be instantiated with different layer types depending on the data modality: $\text{Encoder}(x)=h_{L}$ where $h_{i}=\text{Layer}_{i}(h_{i-1})$ and $h_{0}=x$. Here, $\text{Layer}_{i}$ can be any differentiable layer. The encoder learns condition-invariant features from the source distribution $\mathbb{P}$, while the decoder applies condition-specific transformations. Importantly, this flexibility allows $\text{Layer}_{i}$ to include any standard neural network components.
@@
1:Input: Distributions $\mathbb{P}$, $\mathbb{Q}$, $\mathbb{S}$ accessible by samples; Transport network $T_{\theta}\colon\mathbb{R}^{P}\times\mathbb{R}^{S}\times\mathcal{C}\rightarrow\mathbb{R}^{Q}$; Critic network $f_{\omega}\colon\mathbb{R}^{Q}\times\mathcal{C}\rightarrow\mathbb{R}$; Cost $\mathcal{L}\colon\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$; Conditioning distribution $\mathcal{C}$; Maximum steps NN, Transport iterations per step $K_{T}$
2:Pre-training: Initialize $T_{\theta}$ and $f_{\omega}$ using objectives in Eq. (3)-(7)
3:Output: Learned conditional transport map $T_{\theta}$
4:for $t=1,2,\ldots,N$ do
5:  Sample conditioning $c\sim\mathcal{C}$
6:  Sample batches $Y\sim\mathbb{Q}$, $X\sim\mathbb{P}$; for each $x\in X$ sample $Z_{x}\sim\mathbb{S}$
7:  $\mathcal{L}_{f}\leftarrow\frac{1}{|X|}\sum_{x\in X}\frac{1}{|Z_{x}|}\sum_{z\in Z_{x}}f_{\omega}(T_{\theta}(x,z,c),c)-\frac{1}{|Y|}\sum_{y\in Y}f_{\omega}(y,c)$
8:  Update $\omega$ using $\frac{\partial\mathcal{L}_{f}}{\partial\omega}$ (gradient ascent)
9:  for $k_{T}=1,2,\ldots,K_{T}$ do
10:   Sample conditioning $c\sim\mathcal{C}$
11:   Sample batch $\tilde{X}\sim\mathbb{P}$; for each $x\in\tilde{X}$ sample $\tilde{Z}_{x}\sim\mathbb{S}$
12:   $\mathcal{L}_{T}\leftarrow\frac{1}{|\tilde{X}|}\sum_{x\in\tilde{X}}\left[\mathcal{L}(x,T_{\theta}(x,\tilde{Z}_{x},c))-\frac{1}{|\tilde{Z}_{x}|}\sum_{z\in\tilde{Z}_{x}}f_{\omega}(T_{\theta}(x,z,c),c)\right]$
13:   Update $\theta$ using $\frac{\partial\mathcal{L}_{T}}{\partial\theta}$
14:  end for
@@ 3.3 Conditioning mechanism
Our framework supports conditioning on discrete and continuous variables, with the capability to enable either or both types of conditioning depending on the application. The conditioning module transforms these inputs into a unified representation that modulates the transport map, enabling efficient scaling to applications requiring transport across many conditioning values without solving separate OT problems.
@@
Unified Conditioning. The conditioning module flexibly handles discrete variables, continuous variables, or both. When both types are present, their representations are concatenated, then processed to produce the final conditioning vector $\mathbf{c}=\text{MLP}([\mathbf{e}_{\text{discrete}},\text{PE}(p_{\text{continuous}})])$, where $\cdot$$\cdot$] denotes concatenation and MLP is a multi-layer perceptron with SiLU Ramachandran et al. (2017) activations. The resulting vector $\mathbf{c}\in\mathbb{R}^{d_{c}}$ is used to modulate the transport map in the latent space.
Flexible Configuration. Both discrete and continuous conditioning are optional and can be independently enabled or disabled. The module requires at least one type of conditioning to be active. This flexibility allows our framework to adapt to various application domains—from purely categorical problems to continuous spatiotemporal transport tasks. Unlike CondOT (Bunne et al., 2022), we use different conditioning embeddings for T and f, as this showed better performance. This separation proves to be quantitatively important, suggesting that the transport map and critic function use conditioning information in fundamentally different ways—the transport map must adapt its mapping strategy based on conditions, while the critic must evaluate transport quality conditional on the same information.
@@ 3.4 Conditioning modules variants
[…]
Alternative Conditioning Mechanisms We evaluated several conditioning strategies beyond our hypernetwork approach. The simplest baseline is Concatenation, which directly combines feature vectors with the condition but cannot model the fundamentally different transport behaviors required when conditions significantly alter the optimal mapping strategy. More sophisticated approaches include Cross-Attention (Vaswani et al., 2017), which uses attention mechanisms to modulate features, and Feature-wise Linear Modulation (FiLM) (Perez et al., 2018), which applies learnable transformations to existing features. However, these methods are fundamentally limited to adjusting or re-weighting existing computations rather than enabling entirely different computational paths per condition. We also tested normalization-based methods like Adaptive Instance Normalization (Huang and Belongie, 2017) and Conditional Layer Normalization (Su et al., 2021), along with attention-inspired techniques such as Squeeze-and-Excite (Hu et al., 2018) and Feature-wise Affine Normalization (FAN) (Zhou et al., 2021), but these approaches similarly cannot generate the qualitatively different transport behaviors required when conditions induce fundamentally different optimal mappings. Although these alternatives showed promise in specific scenarios, our hypernetwork approach consistently demonstrated superior performance across all benchmarks. Notably, our lightweight hypernetwork implementation proved to be computationally efficient in training time, making it a sound choice even for low-budget scenarios. We ablate these components on Section 5.
@@ 3.5 Training Procedure
During training, the transport map $T$ and critic function $f$ must maintain adversarial balance while adapting to diverse conditioning values. This creates optimization challenges, where networks must simultaneously learn transport mappings and adapt to diverse conditioning values. We address them through a two-phase approach: pre-training for stable initialization followed by minimax optimization.
[…]
@@
Optimization: Following pre-training, we employ an alternating training that reflects the minimax structure of optimal transport. We detail our approach in Algorithm 1. Following Korotin et al. (2023), we perform $K_{T}>1$ updates for the transport map per critic update. This asymmetry addresses the imbalance in learning complexity—the transport map must learn condition-dependent transformations while maintaining transport constraints, whereas the critic primarily serves as a measure of transport quality. We find $K_{T}\in[4,6]$ provides an adequate balance between training stability and efficiency.This asymmetric training schedule proves essential for conditional transport learning, where the complexity imbalance between transport and critic functions is amplified by the need to handle diverse conditioning values.
@@
Conditions Sampling: The performance of this learning procedure depends on how we sample from the conditioning space $\mathcal{C}$ during training. For discrete variables, we use uniform sampling across all categories. For continuous variables, we employ Beta distribution sampling $\text{Beta}(\alpha,\beta)$ with symmetric parameters ($\alpha=\beta=0.95$), which slightly oversamples values around the min and max values in the training datasets. This strategy prevents underfitting in boundary conditions. This sampling strategy is important for applications like global sensitivity analysis, where transport quality must be maintained across the full range of conditioning values to ensure accurate sensitivity indices.
4 Applications and Experimental Setup
We evaluate our conditional neural transport framework across diverse applications that span scientific computing, sensitivity analysis, and image generation. These applications are specifically chosen to demonstrate the breadth of conditioning scenarios our framework can handle: from purely discrete conditioning to hybrid discrete-continuous conditioning, and from lower-dimensional scientific problems to image generation tasks. Each application presents unique challenges that test different aspects of our conditional transport approach, providing comprehensive validation of the framework’s capabilities.
4.1 Climate Economic Impact Distribution Transport
We examine the economic impacts of climate change using the empirical damage function model of Burke et al. (2015). This application requires building an emulator for complex multivariate distributions conditioned on both categorical (scenario) and continuous (time) variables, representing a challenging test case for hybrid conditioning mechanisms.
Problem Formulation. Let $\mathcal{X}\subset\mathbb{R}^{n}$ represent the space of GDP per capita with climate damages across $n=20$ countries. For each country $i$, the impact is quantified through 1000 bootstrap replicates, resulting in a high-dimensional empirical distribution that captures the uncertainty in damage estimates. We define our conditioning space $\mathcal{C}=\mathcal{C}_{\text{ssp}}\times\mathcal{C}_{\text{year}}$, where $\mathcal{C}_{\text{ssp}}=\{0,1,2,3\}$ represents four SSP scenarios (SSP1-1.9, SSP2-4.5, SSP3-7.0, SSP5-8.5) and $\mathcal{C}_{\text{year}}=[2030,2100]$ represents projection years. Our goal is to learn a conditional transport map $T_{\theta}:\mathbb{R}^{n}\times\mathcal{Z}\times\mathcal{C}\rightarrow\mathcal{X}$ that efficiently transforms samples from a reference distribution to match target distributions under different climate scenarios and future years.
Dataset and Evaluation. The dataset contains empirical damage distributions for 20 countries across 4 SSP scenarios and 71 years (2030-2100), totaling 5,680 conditioning combinations. Each distribution is characterized by 1000 bootstrap samples, providing rich uncertainty quantification. We use 80% of the data for training and 20% for evaluation, ensuring temporal and scenario stratification. We evaluate performance using the Wasserstein-2 distance between generated and target distributions-
4.2 Global Sensitivity Analysis for Integrated Assessment Models
Our second application focuses on global sensitivity analysis (GSA) for the RICE50+ Integrated Assessment Model (IAM) (Gazzotti, 2022), using the optimal transport-based sensitivity indices presented in Borgonovo et al. (2024). This application demonstrates our framework’s ability to replace computationally intensive traditional approaches with efficient neural approximations.
Theoretical Background. Let $\mathbf{Y}=(Y_{1},\dots,Y_{k})$ represent quantities of interest as a function of inputs $\mathbf{C}=(C_{1},\dots,C_{d})$. Assume $\mathbf{C}$ and $\mathbf{Y}$ are random vectors on probability space $(\Omega,\mathcal{B},\mu)$, with $\mu_{C_{i}}$ and $\mu_{\mathbf{Y}}$ denoting the distributions of $C_{i}$ and $\mathbf{Y}$, respectively. Consider a model $\mathbf{f}:\mathcal{C}\subset\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$. The OT-based sensitivity index for the $i$-th input is:
$\iota^{K}(\mathbf{Y},C_{i})=\frac{\mathbb{E}_{C_{i}}[K(\mu_{\mathbf{Y}},\mu_{\mathbf{Y}|C_{i}})]}{\mathbb{E}[k(\mathbf{Y},\mathbf{Y}^{\prime})]},$ (7)
where $K(\mu_{\mathbf{Y}},\mu_{\mathbf{Y}|C_{i}})$ is the OT cost between unconditional and conditional output distributions, and the denominator provides normalization. This index satisfies zero-independence ($\iota^{K}(\mathbf{Y},C_{i})=0$ iff $\mathbf{Y}$ independent of $C_{i}$) and max-functionality ($\iota^{K}(\mathbf{Y},C_{i})=1$ iff $\mathbf{Y}=\mathbf{g}(C_{i})$ for some function $\mathbf{g}$).
Problem Formulation. Let $f:\mathcal{C}\subset\mathbb{R}^{3}\rightarrow\mathcal{Y}\subset\mathbb{R}^{58}$ be the RICE50+ model mapping input parameters $\mathbf{C}\in\mathcal{C}$ to output $\mathbf{Y}\in\mathcal{Y}$. RICE50+ (Gazzotti, 2022) is an IAM with high regional heterogeneity used to assess climate policy benefits and costs. We estimate the sensitivity of CO2 emissions for a single region to three inputs related to emissions abatement costs (klogistic), aversion to inter-country inequality (gamma), and climate impacts (kw_2), leveraging data from Chiani et al. (2025).
Traditional computation of these indices requires solving multiple OT problems for each conditioning variable and partition—specifically, for $d$ variables with $M$ partitions each, this requires $d\times M$ separate OT solutions, becoming computationally intractable for large-scale applications. We partition each input space into $M=25$ bins and define conditioning space $\mathcal{\tilde{C}}=\{0,1,2\}\times[0,1]$, where the first component represents variable index and the second represents normalized partition. Our approach learns a single conditional transport map $T_{\theta}:\mathcal{Y}\times\mathcal{Z}\times\mathcal{\tilde{C}}\rightarrow\mathcal{Y}$ that can compute sensitivity indices for all variables and partitions efficiently.
Dataset and Evaluation. The dataset contains 10,000 model evaluations with corresponding input-output pairs. We use the traditional estimator with partitioned inputs as ground truth, computed using transportation simplex solvers. We evaluate our model with the Pearson correlation $\rho$ between neural and simplex-computed costs.
4.3 Image Generation Tasks
To demonstrate the generality of our approach beyond scientific computing, we evaluate conditional image generation on standard computer vision benchmarks. These tasks test our framework’s ability to handle high-dimensional visual data with both discrete class labels and continuous attribute conditioning.
MNIST Conditional Generation. We perform conditional digit generation with discrete conditioning, using class labels $\mathcal{C}=\{0,\ldots,9\}$. For this application, we use convolutional models for both $T$ and $f$, which include convolutional layers with different values for strides and kernel sizes, therefore leveraging standard components for image generation models without imposing convexity constraints.
Problem Formulation. We learn a transport map $T_{\theta}:\mathcal{Z}\times\mathcal{C}\rightarrow\mathbb{R}^{28\times 28}$ that transforms noise into digit images conditioned on class and attributes.
Dataset and Evaluation. We use the standard MNIST dataset augmented with rotation and stroke thickness variations to create ground truth conditional distributions. To evaluate this setup, we use the Fréchet Inception Distance (FID) for generation quality, and digit classification accuracy.
4.4 Experimental Protocols
Hardware and Implementation. All experiments are conducted on RTX 4070 GPU with i7-13700H CPU. Our framework is implemented in PyTorch with automatic mixed precision training. We use CodeCarbon (Lacoste et al., 2019) to monitor energy consumption across experiments.
Architecture Details. For climate and GSA applications, we use fully connected encoder-decoder architectures with 4 encoder layers (hidden size 128) and 8 decoder layers for transport networks, and 3 encoder/decoder layers for critic networks. For image generation, we employ convolutional architectures with residual blocks. All networks use layer normalization, orthogonal initialization, and learnable residual scaling parameters $\alpha$.
Training Configuration. We employ our two-phase training procedure: 500 steps of pre-training followed by alternating minimax optimization. Transport networks receive $K_{T}=5$ updates per critic update. Learning rates are $2\times 10^{-5}$ for transport and $3\times 10^{-5}$ for critic networks, with AdamW optimization and weight decay of 0.03. Conditioning sampling uses uniform distribution for discrete variables and $\text{Beta}(0.95,0.95)$ for continuous variables.
Experimental Comparisons. Our evaluation focuses on systematic ablation studies to identify optimal design choices within our framework. We compare different conditioning mechanisms including concatenation, cross-attention, Feature-wise Linear Modulation (FiLM), normalization-based approaches (Adaptive Instance Normalization, Conditional Layer Normalization), and attention-inspired techniques (Squeeze-and-Excite, Feature-wise Affine Normalization). We also evaluate different continuous variable encoding strategies (positional encoding, Fourier features, scalar values), architectural choices (shared vs. separate conditioning embeddings), and training procedures (with and without pretraining). For GSA applications, we validate against traditional transportation simplex solvers to establish ground truth sensitivity indices.
Comparisons with CondOT Bunne et al. (2022). We additionally compare against two CondOT baselines. CondOT default uses the original CondOT architecture and optimization recipe, together with a tokenized factorized conditioning scheme in which categorical and discretized continuous condition components are represented as separate conditioning tokens. Since the original CondOT formulation is not designed for mixed categorical: continuous conditioning in the form considered here, continuous variables (time indices or partition levels) must be discretized before being provided to the model. To provide a more informative comparison, we also define a CondOT adapted variant, in which we replace this discretized conditioning with a mixed conditioning mechanism combining one-hot categorical inputs and normalized continuous scalars. We do not include CondOT in the MNIST experiments, since its original formulation is not designed for convolutional architectures. For both configurations, we create architectures with a comparable number of trainable parameters to our models, to fully isolate the impact of architecture design and optimization choices rather than model capacity. Number of training epochs remain equal to our models, while optimization choices (optimizer betas, number of inner iterations) for CondOT follow those of their original implementation.
5 Results and Analysis
We present a comprehensive empirical evaluation of our conditional neural transport framework across the applications described in Section 4. Our evaluation follows a systematic approach: we first establish the effectiveness of our architectural improvements through ablation studies on both unconditional and conditional settings, then demonstrate the practical impact of our approach on real-world applications. Finally, we analyze computational aspects and discuss the implications of our findings.
5.1 Ablation Studies
To validate our design choices and quantify the contribution of each component, we conduct systematic ablation studies examining architectural improvements and conditional-specific design decisions.
5.1.1 Architectural Foundation
We first evaluate architectural improvements to the base NOT framework Korotin et al. (2023), establishing a strong foundation for our conditional extensions. Building upon their ReLU-based architectures, we incrementally add orthogonal initialization, residual connections, and layer normalization, while preserving identical training configuration and comparable parameter counts across models.
Figure 3 presents results across three datasets that test different aspects of neural optimal transport. The first two rows show 2D probability distributions using simple MLPs as baseline architecture. We measure KL divergence and Wasserstein loss between target and transported distributions (from a 2D Uniform prior), using $2^{15}$ samples. The third row showcases a convolutional model performing image-to-Image generation on MNIST (digit 2 to digit 3), evaluated with MMD distance, Wasserstein loss, FID score, and classification accuracy.